please can you plug on calculator and get
1) P (z > -0.091)
and
2) P (z > -0.704)
I look at the z tables but dont know how to find when 3 decimal places). thanks
Answer #1

Answers: 1) P(Z > -0.091) = 1 - P(Z < -0.091) = 1-  0.4637463 =  0.5362537
= 0.536 (rounded to 3 decimal places)

2) P(Z > -0.704) = 1- P(Z < -0.704) = 1 - 0.2407164 =  0.7592836

= 0.759 (rounded to 3 decimal places)

Here Z ~ N(0,1)

The answers are obtained using R-software, using the pnorm() function that gives the cumulative probability distribution of a random variable.

The R-code and output are attached below.

RGui (32-bit) File Edit View Misc Packages Windows Help ALE BEN kara VE R Console > 1-pnorm (-0.091) [3] 0.5362537 1.pnorm(-0